# Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning

### v1

@@ 7 Limitations
Generalization Ability:
While our empirical investigation has demonstrated promising results for the Mistral-7b-instruct model, additional benchmark evaluations may be needed to validate the generalization ability of the proposed Smurfs framework across various LLM architectures and tasks.
@@
Due to device limitations and computational constraints, our experiments primarily focused on the 7B models. Further evaluations with larger and smaller LLMs are required to assess the impact of the Smurfs framework on models of different sizes.
Computational Efficiency:
Although preliminary findings suggest that the Smurfs framework can enhance computational efficiency through efficient verification, a more detailed analysis is needed to quantify the computational overhead introduced by the multi-agent architecture.
@@
Appendix A Pilot Study Settings
Figure 5: Demonstration of the memory of the Smurfs framework.
A.1 Pilot Study 1
We conducted the first pilot study using the ToolBench benchmark Qin et al. (2024), which comprises over 16,000 APIs focused on tool usage and planning. We divided the benchmark into two subsets: "ToolBench-long," containing samples requiring more than three steps for ChatGPT-DFSDT, and "ToolBench-all," which includes all samples from ToolBench. On average, we consider the "ToolBench-long" subset to be more complex than the "ToolBench-all" set. We maintained all other settings consistent with the original source settings.
A.2 Pilot Study 2
The second pilot study is conducted using the BMTools Qin et al. (2023) test datasets. We augmented each sample by adding random, irrelevant test samples as prefix prompts to increase their input length. We used tool selection accuracy as the metric to evaluate the LLMs’ model selection capabilities. The experiment uses Mistral-7B-Instruct-v0.1 Jiang et al. (2023) and Qwen1.5-7B-Chat Bai et al. (2023) for evaluation.
@@ Appendix B Details of the Smurfs
B.1 Memory management of Smurfs
When each agent performs a task, the prompt information they receive varies. We refer to this as the memory management system of Smurfs. As illustrated in Figure 5, information is primarily divided into four categories. The first category is local memory, consisting of the thought and answer from the previous steps. When backtracking happens, the thought and answer of the backtracking step will be popped out from the local memory. The second category is global memory, which stores all the previous steps history including those that have been backtracked. The third category is the tool list, similar to other tool-using frameworks, storing brief information about all accessible tools. Lastly, the tool document provides detailed usage information like the parameters infomation for each tool listed.
As illustrated in Figure 3, each agent has its own memory context. The Executor Agent’s memory will be discussed in the next section. The Answer Agent takes in the subtask, the tool response and the local memory to generate the answers to the subtask using all the local memory so far. Then the thought and answer of this step will be added to the local memory and the global memory. The Verifier Agent takes in the subtask, the answer from the Answer Agent and return the status of the subtask. If the status is solved, then the system will return the subtask’s answer and process to the next subtask; If the status is pending, the Verifier Agent will give hint to the Executor Agent to process to the next step. When the max steps have been reached for a subtask, the answer agent will generate the answer for the subtask using the global memory and process to the next subtask.
B.2 Executor Agent Workflow
Next, let’s illustrate how memory is accessed during the Executor Agent’s working process, as shown in Figure 6. When a subtask arises, the Executor Agent first receives the hint (task instruction) from the Verifier Agent, the subtask, the available tool list to generate the thought(a strategy for task execution) at this step. It then takes in the subtask, the thought and the tool list to generate an action with a specific action name. Subsequently, it takes in the subtask, the thought, the tool document of the chosen action and the local memory to generate the input of the chosen action.
B.3 Backtracking Mechanism
Smurfs uses a backtracking mechanism similar to DFSDT  Qin et al. (2024). As illustrated in Figure 3, when tool responses error, the system will delete the tool from the tool list and retry this step. However, when all tools have reported error or the model thinks the available tools can’t solve the subtask at this step, the system will pop out the latest local memory and the latest step, delete the tool used by the popped out step from the tool list of the current step and retry the the other solution path.
Appendix C Prompts for multi-agent implementation

### v3

@@ 7 Limitations
Due to computational constraints, our experiments did not include larger and more diverse types of LLMs.
Agent Component Scale-Up:
Although we selected the most common and intuitive agent roles for the proposed MAS, there are many possibilities for researchers to explore. Investigating more well-designed agent roles may help improve the effectiveness of the agent system, and developing automated methods to identify these roles could facilitate effective scaling.
@@
Appendix A Details of DFSDT
See Figure 5.
Figure 5: A comparison of DFSDT and conventional CoT or ReACT during model reasoning (left) Qin et al. (2024). We show part of the solution path annotation process using GPT-4 (right).
@@ Appendix B Details of the Smurfs
See Figure 6 for executor working process and Figure 4 for memory and tool library of Smurfs.
@@
Appendix C Experiment Settings for Hotpot QA
Following settings in Qiao et al. (2024), which is randomly select 300 dev questions divided into three levels for evaluation, with 100 questions in each level. For tool library that can be used in HotpotQA, see Table 6
Name Definition Usage
  BingSearch BingSearch engine can search for rich knowledge on the internet based on keywords, which can compensate for knowledge fallacy and knowledge outdated. BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query. Be specific and precise with your query to increase the chances of getting relevant results. For example, Bingsearch[popular dog breeds in the United States]
Retrieve Retrieve additional background knowledge crucial for tackling complex problems. It is especially beneficial for specialized domains like science and mathematics, providing context for the task Retrieve[entity], which retrieves the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to retrieve. For example, Retrieve[Milhouse]
Lookup A Lookup Tool returns the next sentence containing the target string in the page from the search tool, simulating Ctrl+F functionality on the browser. Lookup[keyword], which returns the next sentence containing the keyword in the last passage successfully found by Retrieve or BingSearch. For example, Lookup[river].
Table 6: Tool library for HotpotQA.
Appendix D Prompts for multi-agent implementation
@@
Appendix E Token Cost on StableToolBench Evaluation
We analyzed the token cost for the StableToolBench experiments. As shown in Table 7, the total token cost for each subtask within the StableToolBench is compared across three candidate tool-planning methods. The data demonstrates that, across all tasks from easy to hard, DFSDT consistently incurs high token costs, while the other two methods maintain relatively low token costs. This verifies the context-efficiency of the proposed method.
Backbone Method StableToolBench
I1-Inst. I1-Cat. I1-Tool. I2-Cat. I2-Inst. I3-Inst. Average
Total Avg. Total Avg. Total Avg. Total Avg. Total Avg. Total Avg. Total Avg.
GPT-3.5 Turbo ReACT 1,010,304 6,198 824,676 5,390 1,010,514 6,396 900,855 7,265 824,510 7,778 461,121 7,559 838,663 6,764
GPT-3.5 Turbo DFSDT 3,303,062 20,264 2,745,667 17,945 3,152,532 19,953 2,560,297 20,648 3,098,365 29,230 1,390,787 22,800 2,708,452 21,807
GPT-3.5 Turbo Smurfs 1,090,404 7,127 1,917,348 11,763 1,464,535 9,269 957,088 7,638 1,096,162 10,341 632,084 10,362 1,191,270 9,417
Table 7: Token costs for various candidate tool-planning methods on the StableToolBench benchmark Guo et al. (2024). ‘Total’ indicates the total number of tokens used to complete each subtask, including both prompt and completion tokens. ‘Avg.’ represents the average number of tokens used per question within the subtasks. Higher token counts imply greater costs for solving the same task.